# Dynamically localized but deconfined excitations in strained classical spin ice

### v1

@@ Abstract
We study classical spin ice under uniaxial strain along the $[111]$ crystallographic axis. Remarkably, such strain preserves the extensive ice degeneracy and the corresponding classical Coulomb phase. The emergent monopole excitations remain thermodynamically deconfined exactly as in the isotropic case. However, their motion under local heat bath dynamics depends qualitatively on the sign of the strain. In the low-temperature limit for negative strain, the monopoles diffuse, while for positive strain, they localize. Introducing additional ring exchange dynamics into the ice background transforms the localized monopoles into sub-dimensional excitations whose motion is restricted to diffusion in the (111)-plane. The phenomena we identify are experimentally accessible in rare-earth pyrochlores under uniaxial pressure as well as in tripod kagome materials. The diffusive versus localized nature of the monopoles manifests in characteristic magnetic noise spectra, which we compute.
@@ I Introduction
The nearest-neighbor Ising model on the pyrochlore lattice, called spin ice, is the quintessential frustrated magnetic system [1, 2, 3]. Its unusual properties include extensive zero-point entropy [4] and fractionalized excitations in the form of thermodynamically deconfined magnetic monopoles [5, 6]. This work focuses on spin ice under uniaxial $[111]$-strain, see Fig. 1. Such strain preserves the extensive entropy corresponding to the thermodynamic Coulomb phase. Nonetheless, the strain leads to dramatic changes in the dynamics of the monopoles.
(a) Uniaxial pressure.
@@
Figure 1: Two illustrations of the [111]-deformed pyrochlore with symmetry-reduced space group $R\bar{3}m$ describing the nearest-neighbor spin ice model in Eq. (missing) 1. Distorted bonds along the $[111]$ direction are highlighted with dashed lines. (a) Strain induced by uniaxial pressure where bonds along the $[111]$ direction are squeezed, increasing the interaction strength. (b) Substitution of the atoms located on the $[111]$-sublattice (green) results in a symmetry-equivalent modulation of the nearest-neighbor interaction. Tripod kagome materials [7, 8] realize the extreme case where the green atoms are non-magnetic and the kagome planes decouple.
@@
In the minimal, “standard model” of spin ice dynamics, single spins flip stochastically according to a Markovian heat-bath [9, 10]. This leads to monopole diffusion in the isotropic case. Remarkably, the introduction of $[111]$-strain leads to qualitatively distinct monopole dynamics depending on the sign of the strain. If the strain increases the interactions between spins separated in the $[111]$-direction, the monopoles completely localize. In the opposite scenario, monopoles remain diffusive with a weak anisotropic correction to the diffusion tensor. Both, the dynamical localization and the anisotropic diffusion, manifest in macroscopic dynamical properties such as the magnetic noise spectrum [11, 12], which has been of much recent experimental interest. The zero-temperature thermodynamics are completely insensitive to (moderate amounts of) strain; the monopoles are thus excitations that are thermodynamically deconfined but dynamically localized.
@@
If one additionally includes stochastic ring exchanges of the ice background on top of the standard model dynamics, the localized monopoles become diffusive in the $(111)$-plane, leading to sub-dimensional diffusion. In this sense, they are classical “sub-dimensional excitations” [13, 14].
@@
Figure 2: The phase diagram of strained spin ice, $H_{S}$, as a function of distortion, $\delta$, and temperature, $T\,(J)$. The model realizes two ordered and two disordered low-temperature phases that are connected by the duality axis at $\delta=-1$. For $\delta<-3$, the model is long-range ordered following an all-in-all-out pattern (yellow). Under duality, Eq. (missing) 4, the ordered state maps to a ferromagnetic state at $\delta>2$ (green). Similarly, the duality maps the famous spin ice phase (blue) for $-1<\delta<1$ to an equivalent disordered Coulomb phase (red), we name inverted spin ice for $-3<\delta<-1$. The model changes character at $\delta=-1$, the duality axis, which marks a transition from ferro- to anti-ferromagnetic couplings along $[111]$. At the duality point, $[111]$ spins decouple, and the system is composed of alternating anti-ferromagnetic kagome planes sandwiched between planes of free spins. The white-shaded regions indicate the temperature regimes where the dynamics are strongly influenced by strain; in the A regimes, monopoles are (anisotropically) diffusive, whereas they are dynamically localized in the B regimes. Again, the duality provides the analogous dynamical description in the inverted spin ice phase.
@@
Our model is experimentally motivated, and moderate values of strain can be induced by pressure along the $[111]$ crystallographic axis, see 1(a). Experiments on the classical spin ice compounds Dy2Ti2O7 and Ho2Ti2O7 have been carried out with uniaxial pressure applied along the $[100]$, $[110]$, and $[111]$ axes [15, 16, 17, 18]. These experiments probed static quantities, which are insensitive to the monopole localization we predict.
@@
Atomic substitution is an alternate route to exploring spin ice with the same reduced symmetry as the $[111]$-strained pyrochlore. By systematically substituting atoms on a particular sublattice, one obtains larger effective strain parameters, see 1(b). This could even flip the sign of the interactions between different species relative to the usual spin ice. A special case of such materials is provided by kagome Ising magnets such as Dy3Mg2Sb2O14 [7, 8]. These realize the extreme case where one sublattice spin is non-magnetic, leaving effectively decoupled kagome layers.
@@
Thermodynamic properties of spin ice under uniaxial strain have been investigated theoretically in the past [19, 20]. However, the main focus of previous works was the effect of pressure applied along the $[100]$ direction, where any amount of strain breaks the extensive degeneracy of the ice states, resulting in qualitatively different behavior compared to the $[111]$ case we consider.
@@ II Ground states and excitations
The first sum runs over pairs with one spin on sublattice $\mathcal{S}_{[111]}$ – the sublattice on which spins align with the $[111]$ direction – and the second sum runs over pairs with no spin on sublattice $\mathcal{S}_{[111]}$. We denote the other three sublattices collectively as $\mathcal{S}_{K}$ as these spins form kagome planes. We take $J>0$ as the interaction strength. In the $\sigma=\pm 1$ pseudospin Ising convention used in Eq. (missing) 1, $J>0$ corresponds to anti-ferromagnetic interactions. In the vector spin language used to describe spin ice materials like Dy2Ti2O7 and Ho2Ti2O7, which we use pictorially throughout the manuscript, the spin-spin interaction is, in fact, ferromagnetic [3]. $\delta$ is a dimensionless scalar parameter that modulates the interaction strength and can take both positive and negative values. A non-zero $\delta$ reduces the space group of the lattice from the usual $Fd\bar{3}m$ of the pyrochlore lattice to $R\bar{3}m$ [21, 22, 23, 8]. As in conventional spin ice, we can re-express $H_{\rm nn}$ up to a constant term as a sum over all tetrahedra,
@@
where the spins $\sigma_{t,1}$ are members of $\mathcal{S}_{[111]}$.
@@
Figure 3: Left: All 16 spin configurations on a single tetrahedron organized into four degenerate groups according to $H_{S}$, Eq. (missing) 2. The energies of the groups depend on the strain parameter, $\delta$, and each phase in Fig. 2 corresponds to the regime where energy is minimized by tiling the system with configurations from a particular group. Type 0 and type 1A configurations, respectively, correspond to the spin ice and inverted spin ice phases, with an extensive number of possible ground states. In contrast, the ferromagnetic (type 1B) and the all-in-all-out (type 2) phases are long-range ordered with only two ground state configurations each. The duality transformations Eq. (missing) 4 is a unique mapping between type 0 and type 1A as well as type 1B and type 2 configurations. Monopoles and anti-monopoles are highlighted in blue and red, respectively. The sublattice $\mathcal{S}_{[111]}$ spin is located on the top of the tetrahedra and couples to the three $\mathcal{S}_{K}$ spins with interaction strength $J(1+\delta)$, illustrated by dashed bonds. The three $\mathcal{S}_{K}$ spins interact with strength $J$, visualized by solid lines. Right: Energy versus $\delta$ for the four configuration groups. The colors of the lines and background correspond to the phases in Fig. 2. It further highlights the most relevant excitation type at each value of $\delta$. For example, within the spin ice phase, the type 1A monopoles (red) dominate for $-1<\delta<0$, and type 1B monopoles (green) dominate for $10<\delta<$ 1.
@@
Much of the model can be understood by examining the $162^{4}=$ 16 spin configurations on a single tetrahedron, see Fig. 3. Six of these have two spins pointing in and two spins pointing out of the tetrahedron; we refer to these states as Type 0 configurations. Pointing “in” and “out” corresponds to $\sigma=+1$ and $\sigma=-1$ in the notation of Eq. (missing) 2. Any configuration where every tetrahedron follows this “two-in-two-out” rule, also known as the “ice rule” [24], is a ground state of isotropic nearest-neighbor spin ice. As pointed out by Anderson [1], there is an extensive entropy of such ground states, originally estimated by Pauling [25] in the context of water ice [26]. Flipping a single spin out of a two-in-two-out ground state creates a pair of excitations that take the form of three-in-one-out and one-in-three-out tetrahedra. These are the famous emergent magnetic monopoles [5], which we call Type 1 configurations. Once created, monopoles can move independently through further spin flips and only annihilate if they meet a monopole of the opposite sign. Eight of the sixteen spin configurations on a tetrahedron correspond to a single monopole. The remaining two “double monopole” configurations have all four spins pointing in the same direction, and we call these Type 2 configurations.
@@
While type 0 and type 2 configurations remain degenerated, the $[111]$-strain lifts the degeneracy of the eight type 1 configurations. The single monopoles are type 1A with energy $J(2+\delta)^{2}/2$ if the minority spin is not on $\mathcal{S}_{[111]}$, and type 1B with energy $J(2-\delta)^{2}/2$ if the minority spin is on $\mathcal{S}_{[111]}$. The minority spin is the out spin in a three-in-one-out tetrahedron and the in spin in a three-out-one-in tetrahedron. The two type 1B configurations with minority spin on sublattice $\mathcal{S}_{[111]}$ are energetically “cheaper” for $\delta>0$, and monopoles with the minority spin on any of the other sublattices are favored for $\delta<0$. Monopoles can change type when they move, but this carries an associated energy cost. Dynamical localization of monopoles occurs when the temperature is low compared to the cost of converting a type 1B to a type 1A monopole, as further explained in Sec. IV.
@@
For $\delta>1$, the energy of the type 1B monopoles $E_{1B}=J(2-\delta)^{2}/2$ is lower than the energy of any other group. Ground states are formed by a complete tiling of type 1B monopoles on every tetrahedron with all minority spins on sublattice $\mathcal{S}_{[111]}$. There are only two configurations that minimize the energy on all tetrahedra simultaneously, corresponding to polarizing all spins along or against the $[111]$ direction. The system undergoes a phase transition into this spontaneous symmetry-broken ferromagnetic state, where the transition temperature $T_{c}$ grows with increasing $\delta$. Phrased differently, this is a pressure-induced crystallization of a cooperative paramagnet [27]. The ferromagnetic phase persists up to arbitrarily large $\delta$. The lowest-energy excitations in this regime are two-in-two-out (type 0) states. A single spin flip out of the ground state can create a pair of such defects, but they are not mobile – there are no deconfined excitations in the ferromagnetic phase.
@@
For $-3<\delta<-1$, the energy is minimized by placing a single type 1A monopole on each tetrahedron. Type 1A configurations have a majority spin on sublattice $\mathcal{S}_{[111]}$, resulting in much more freedom in how to form ground states compared to the type 1B tiled ferromagnetic phase. In fact, there is a direct one-to-one mapping between the monopole-favoring ground states for $-3<\delta<-1$ and the two-in-two-out ground states for $-1<\delta<1$ due to the duality around $\delta=-1$. Any ice configuration can be mapped uniquely by the duality transformation, Eq. (missing) 4, to a ground state for $-3<\delta<-1$, and we dub this phase inverted spin ice.
@@
For $\delta<-3$, the double monopole (type 2) energy is the lowest of the four groups and the system orders into an all-in-all-out state. There are two such states with a broken tetrahedral sublattice symmetry, where all spins pointing either from down to up tetrahedra or from up to down tetrahedra. In the language of Ising spin, this refers to all spins being $1$ or $-1$. This ordered phase persists for arbitrarily negative $\delta$, and again leads to a symmetry-breaking phase transition at a critical temperature $T_{c}$ that grows with decreasing $\delta<-3$. As before, the duality transformation maps the all-in-all-out states to the two ferromagnetically ordered ground states at $\delta>1$.
Figure 4: The combined thermodynamic data for $H_{S}$ in Eq. (missing) 2 with $\delta\geq-1$. The left and central columns cover $\delta$ values that correspond to the spin ice regime, whereas the right column refers to the ferromagnetic region at $\delta>1$. Top row: Panels (a,b,c) display the specific heat with the characteristic Schottky anomalies associated with different excitations shown in Fig. 3. Within the spin ice phase, panels (a,b), the two distinct peaks are linked to non-degenerate monopole excitations, which collapse into a single smooth peak for degeneracies occurring at $\delta=0,\pm 1$. In the ferromagnetic regime, the phase transition manifests as a sharp peak marked by dotted vertical lines. Middle row: Single monopole densities, including both type 1A and 1B, but not counting double monopoles. As the temperature is reduced, the monopole density goes to zero in the spin ice regime, panels (d,e), and to one in the ferromagnetic regime, panel (f). For the special points, $\delta=\pm 1$, single monopoles and ice states share the same energy, and the monopole density goes to $1/2$ and $1/4$, respectively. Bottom row: The entropy per spin is shown for the spin ice regime (g,h), matching Pauling estimates as described. Panel (i) of this row displays the static susceptibility measured in the $[111]$ direction for $\delta>1$, showing sharp peaks due to the phase transition.

### v2

@@ Abstract
We study classical spin ice under uniaxial strain along the $[111]$ crystallographic axis. Remarkably, such strain preserves the extensive ice degeneracy and the corresponding classical Coulomb phase. The emergent monopole excitations remain thermodynamically deconfined exactly as in the isotropic case. However, their motion under local heat bath dynamics depends qualitatively on the sign of the strain. In the low-temperature limit for negative strain, the monopoles diffuse, while for positive strain, they localize. Introducing additional ring exchange dynamics into the ice background transforms the localized monopoles into subdimensional excitations whose motion is restricted to diffusion in the (111)-plane. The phenomena we identify are experimentally accessible in rare-earth pyrochlores under uniaxial stress as well as in tripod kagome materials. The diffusive versus localized nature of the monopoles manifests in characteristic magnetic noise spectra, which we compute.
@@ I Introduction
The nearest-neighbor Ising model on the pyrochlore lattice, called spin ice, is the quintessential three-dimensional frustrated magnetic system [1, 2, 3]. Its unusual properties include extensive zero-point entropy [1, 4] and fractionalized excitations in the form of thermodynamically deconfined magnetic monopoles [5, 6]. This work focuses on spin ice under uniaxial $[111]$-strain; see Fig. 1. Such strain preserves the extensive entropy corresponding to the thermodynamic Coulomb phase. Nonetheless, the strain leads to dramatic changes in the dynamics of the monopoles.
(a) Uniaxial stress.
@@
Figure 1: Two illustrations of the [111]-deformed pyrochlore with symmetry-reduced space group $R\bar{3}m$ describing the nearest-neighbor spin ice model in Eq. (missing) 2. Distorted bonds along the $[111]$ direction are highlighted with dashed lines. (a) Strain induced by uniaxial stress where bonds along the $[111]$ direction are squeezed, increasing the interaction strength. (b) Substitution of the atoms located on the $[111]$-sublattice (green) results in a symmetry-equivalent modulation of the nearest-neighbor interaction. Tripod kagome systems [7, 8] realize the extreme case where the green atoms are nonmagnetic and the kagome planes decouple.
@@
In the minimal, “standard model” of spin ice dynamics, single spins flip stochastically according to a Markovian heat-bath [9], which leads to monopole diffusion in the isotropic case [10]. Remarkably, the introduction of $[111]$-strain leads to qualitatively distinct monopole dynamics depending on the sign of the strain. If the strain increases the interactions between spins separated in the $[111]$-direction, the monopoles completely localize. In the opposite scenario, monopoles remain diffusive with a weak anisotropic correction to the diffusion tensor. Both the dynamical localization and the anisotropic diffusion manifest in macroscopic dynamical properties such as the magnetic noise spectrum [11, 12], which has been of much recent experimental interest. Moderate amounts of strain preserve the ice degeneracy of all two-in-two-out states, and the monopoles remain thermodynamically deconfined excitations but become dynamically localized. By this, we mean that a monopole excitation can be moved any distance at a finite, or even zero, energy cost; however, in the low-temperature limit, the timescale on which a monopole moves even a short distance diverges.
@@
If one additionally includes fluctuations in the ice background on top of the standard model dynamics, then the localized monopoles become diffusive in the $(111)$-plane, leading to subdimensional diffusion. In this sense, they are classical “subdimensional excitations” [13, 14].
@@
Figure 2: The phase diagram of strained spin ice, described by the Hamiltonian $H_{nn}$ (Eq. (missing) 2), as a function of distortion, $\delta$, and temperature, $T\,(J)$. The model realizes two ordered and two disordered low-temperature phases that are connected by the duality axis at $\delta=-1$. For $\delta<-3$, the model is long-range ordered following an all-in-all-out pattern (yellow). Under duality, Eq. (missing) 5, the ordered state maps to a ferromagnetic state at $\delta>2$ (green). Similarly, the duality maps the famous spin ice phase (blue) for $-1<\delta<1$ to an equivalent disordered Coulomb phase (red) we name inverted spin ice for $-3<\delta<-1$. The model changes character at $\delta=-1$, the duality axis, which marks a transition from ferro- to antiferromagnetic couplings along $[111]$. At the duality point, $[111]$ spins decouple, and the system is composed of alternating antiferromagnetic kagome planes sandwiched between planes of free spins. The white-shaded regions indicate the temperature regimes where the dynamics are strongly influenced by strain; in the A regimes, monopoles are (anisotropically) diffusive, whereas they are dynamically localized in the B regimes. Again, the duality provides the analogous dynamical description in the inverted spin ice phase.
@@
Our model is experimentally motivated, and moderate values of strain can be induced by stress along the $[111]$ crystallographic axis, see 1(a). Experiments on the classical spin ice compounds Dy2Ti2O7 and Ho2Ti2O7 have been carried out with uniaxial stress applied along the $[100]$, $[110]$, and $[111]$ axes [15, 16, 17, 18]. These experiments probed static quantities, which are insensitive to the monopole localization we predict.
@@
Atomic substitution is an alternate route to exploring spin ice with the same reduced symmetry as the $[111]$-strained pyrochlore. By systematically substituting atoms on a particular sublattice, one obtains larger effective strain parameters, see 1(b). This could even flip the sign of the interactions between different species relative to the usual spin ice. A special case of such materials is provided by kagome Ising magnets such as Dy3Mg2Sb2O14 [7, 8]. These realize the extreme case where one sublattice spin is nonmagnetic, leaving effectively decoupled kagome layers.
@@
Thermodynamic properties of spin ice under uniaxial strain have been investigated theoretically in the past [19, 20]. However, the main focus of previous works was the effect of stress applied along the $[100]$ direction, where any amount of strain breaks the extensive degeneracy of the ice states, resulting in qualitatively different behavior compared to the $[111]$ case we consider.
@@ II Ground states and excitations
$H_{\rm nn}=-3J(1+\delta)\sum_{\mathclap{\begin{subarray}{c}\langle i,j\rangle\\ i\in\mathcal{S}_{[111]},\ j\notin\mathcal{S}_{[111]}\end{subarray}}}\vec{s}_{i}\cdot\vec{s}_{j}-J\sum_{\mathclap{\begin{subarray}{c}\langle i,j\rangle\\ i,j\notin\mathcal{S}_{[111]}\end{subarray}}}\vec{s}_{i}\cdot\vec{s}_{j}\,.$ (1)
where $\delta$ is a dimensionless scalar parameter that modulates the interaction strength and can take both positive and negative values. The first sum runs over pairs with one spin on sublattice $\mathcal{S}_{[111]}$ – the sublattice on which spins align with the $[111]$ direction – and the second sum runs over pairs with no spin on sublattice $\mathcal{S}_{[111]}$. We denote the other three sublattices collectively as $\mathcal{S}_{k}$ as these spins form kagome planes. The variables $\vec{s}$ are classical vector spins forced to point along the local $\langle 111\rangle$ axes as illustrated in  Fig. 3. These can be expressed in terms of Ising variables $\sigma_{i}=\pm 1$ as $\vec{s}_{i}=\sigma_{i}\vec{e}_{i}$, where $\vec{e}_{i}$ is the corresponding unit vector, allowing us to rewrite the Hamiltonian as (see App. A)
@@
We take $J>0$ as the (ferromagnetic) interaction strength, which corresponds to an antiferromagnetic interaction in the language of the Ising variables. A nonzero $\delta$ reduces the space group of the lattice from the usual $Fd\bar{3}m$ of the pyrochlore lattice to $R\bar{3}m$ [21, 22, 23, 8]. As in conventional spin ice, we can re-express $H_{\rm nn}$ up to a constant term as a sum over all tetrahedra,
@@
where $\sigma_{t,1}$ are members of $\mathcal{S}_{[111]}$.
@@
Figure 3: Left: All 16 spin configurations on a single tetrahedron organized into four degenerate groups according to $H_{S}$, Eq. (missing) 3. The energies of the groups depend on the strain parameter, $\delta$, and each phase in Fig. 2 corresponds to the regime where energy is minimized by tiling the system with configurations from a particular group. Type 0 and type 1A configurations, respectively, correspond to the spin ice and inverted spin ice phases, with an extensive number of possible ground states. In contrast, the ferromagnetic (type 1B) and the all-in-all-out (type 2) phases are long-range ordered with only two ground state configurations each. The duality transformations Eq. (missing) 5 is a unique mapping between type 0 and type 1A as well as type 1B and type 2 configurations. Monopoles and anti-monopoles are highlighted in blue and red, respectively. The sublattice $\mathcal{S}_{[111]}$ spin is located on the top of the tetrahedra and couples to the three $\mathcal{S}_{k}$ spins with interaction strength $J(1+\delta)$, illustrated by dashed bonds. The three $\mathcal{S}_{k}$ spins interact with strength $J$, visualized by solid lines. Right: Energy versus $\delta$ for the four configuration groups. The colors of the lines and background correspond to the phases in Fig. 2. It further highlights the most relevant excitation type at each value of $\delta$. For example, within the spin ice phase, the type 1A monopoles (red) dominate for $-1<\delta<0$, and type 1B monopoles (green) dominate for $10<\delta<$ 1.
@@
Much of the model can be understood by examining the $162^{4}=$ 16 spin configurations on a single tetrahedron, see Fig. 3. Six of these have two spins pointing in and two spins pointing out of the tetrahedron; we refer to these states as Type 0 configurations. Any configuration where every tetrahedron follows this “two-in-two-out” rule, also known as the “ice rule” [24], is a ground state of isotropic nearest-neighbor spin ice. As pointed out by Anderson [1], there is an extensive entropy of such ground states, originally estimated by Pauling [25] in the context of water ice [26]. Flipping a single spin out of a two-in-two-out ground state creates a pair of excitations that take the form of three-in-one-out and one-in-three-out tetrahedra. These are the famous emergent magnetic monopoles [5], which we call Type 1 configurations. Once created, monopoles can move independently through further spin flips and only annihilate if they meet a monopole of the opposite sign. Eight of the sixteen spin configurations on a tetrahedron correspond to a single monopole. The remaining two “double monopole” configurations have all four spins pointing either in or out, and we call these Type 2 configurations.
@@
While type 0 and type 2 configurations remain degenerate, the $[111]$-strain lifts the degeneracy of the eight type 1 configurations. The single monopoles are type 1A with energy $J(2+\delta)^{2}/2$ if the minority spin is not on $\mathcal{S}_{[111]}$, and type 1B with energy $J(2-\delta)^{2}/2$ if the minority spin is on $\mathcal{S}_{[111]}$. The minority spin is the out spin in a three-in-one-out tetrahedron and the in spin in a three-out-one-in tetrahedron. The two type 1B configurations with minority spin on sublattice $\mathcal{S}_{[111]}$ are energetically “cheaper” for $\delta>0$, and monopoles with the minority spin on any of the other sublattices are favored for $\delta<0$. Monopoles can change type when they move, but this carries an associated energy cost. Dynamical localization of monopoles occurs when the temperature is low compared to the cost of converting a type 1B to a type 1A monopole, as further explained in Sec. IV.
@@
For $\delta>1$, the energy of the type 1B monopoles $E_{1B}=J(2-\delta)^{2}/2$ is lower than the energy of any other group. Ground states are formed by a complete tiling of type 1B monopoles on every tetrahedron with all minority spins on sublattice $\mathcal{S}_{[111]}$. There are only two configurations that minimize the energy on all tetrahedra simultaneously, corresponding to polarizing all spins along or against the $[111]$ direction. The system undergoes a phase transition into this spontaneous symmetry-broken ferromagnetic state, where the transition temperature $T_{c}$ grows with increasing $\delta$. Phrased differently, this is a stress-induced crystallization of a cooperative paramagnet, akin to phenomena observed in Tb2Ti2O7 [27]. The ferromagnetic phase persists up to arbitrarily large $\delta$. The lowest-energy excitations in this regime are two-in-two-out (type 0) states. A single spin flip out of the ground state can create a pair of such defects, but they are not mobile – there are no deconfined excitations in the ferromagnetic phase.
@@
For $-3<\delta<-1$, the energy is minimized by placing a single type 1A monopole on each tetrahedron. Type 1A configurations have a majority spin on sublattice $\mathcal{S}_{[111]}$, resulting in much more freedom in how to form ground states compared to the type 1B tiled ferromagnetic phase. In fact, there is a direct one-to-one mapping between the monopole-favoring ground states for $-3<\delta<-1$ and the two-in-two-out ground states for $-1<\delta<1$ due to the duality around $\delta=-1$. Any ice configuration can be mapped uniquely by the duality transformation, Eq. (missing) 5, to a ground state for $-3<\delta<-1$, and we dub this phase inverted spin ice.
@@
For $\delta<-3$, the double monopole (type 2) energy is the lowest of the four groups and the system orders into an all-in-all-out state. There are two such states with a broken tetrahedral sublattice symmetry, where all spins pointing either from down to up tetrahedra or from up to down tetrahedra. In the language of Ising variables, this refers to all $\sigma_{i}$ being $+1$ or $-1$. This ordered phase persists for arbitrarily negative $\delta$, and again leads to a symmetry-breaking phase transition at a critical temperature $T_{c}$ that grows with decreasing $\delta<-3$. As before, the duality transformation maps the all-in-all-out states to the two ferromagnetically ordered ground states at $\delta>1$.
Figure 4: The combined thermodynamic data for $H_{S}$ in Eq. (missing) 3 with $\delta\geq-1$. The left and central columns cover $\delta$ values that correspond to the spin ice regime, whereas the right column refers to the ferromagnetic region at $\delta>1$. Top row: Panels (a,b,c) display the specific heat with the characteristic Schottky anomalies associated with different excitations shown in Fig. 3. Within the spin ice phase, panels (a,b), the two distinct peaks are linked to nondegenerate monopole excitations, which collapse into a single smooth peak for degeneracies occurring at $\delta=0,\pm 1$. In the ferromagnetic regime, the phase transition manifests as a sharp peak marked by dotted vertical lines. Middle row: Single monopole densities, including both type 1A and 1B, but not counting double monopoles. As the temperature is reduced, the monopole density goes to zero in the spin ice regime, panels (d,e), and to one in the ferromagnetic regime, panel (f). For the special points, $\delta=\pm 1$, single monopoles and ice states share the same energy, and the monopole density goes to $1/2$ and $1/4$, respectively. Bottom row: The entropy per spin is shown for the spin ice regime (g,h), matching Pauling estimates as described. Panel (i) of this row displays the static susceptibility measured in the $[111]$ direction for $\delta>1$, showing sharp peaks due to the phase transition.